OpenStudy (anonymous):

Find the y-intercept and the x-intercept for the equation 5x+2y=10

OpenStudy (anonymous):

for y intercept set x=0 for x intercept set y=0

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Find the slope if it exists x+3y=6

OpenStudy (anonymous):

isolate y

OpenStudy (anonymous):

What is the y intercept y=1/3x

OpenStudy (anonymous):

*y=-1/3 x +2

OpenStudy (anonymous):

x+3y≤9 3x+y≤9 x≥0 y≥0